OpenStudy (anonymous):

the slope of the line passing through (0,-2) and (3,4)

OpenStudy (anonymous):

2/3 is the slope

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

hero (hero):

-2 = 0m + b -2 = b 4 = 3m + b 4 = 3m - 2 4 + 2 = 3m 6 = 3m 6/3 = m 2 = m The slope is 2